OpenStudy (anonymous):

how do I solve 2^x = 4^x + 1

OpenStudy (anonymous):

2^x = 2^2x + 1 2^x = a a = a^2 +1 a^2 - a + 1

OpenStudy (anonymous):

a^2 - a + 1 = 0 Solve for "a" first

OpenStudy (anonymous):

a= -a^2-1 ?

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

sorry is it -2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=a%5E2+-+a+%2B+1

OpenStudy (anonymous):

thank you